### v2

# Evelyn Hubbard, Liam Cregg, Serdar Yüksel

Abstract
We study the problem of joint optimization involving coding and control policies for a controlled Markovian sytem over a finite-rate noiseless communication channel. While structural results on the optimal encoding and control have been obtained in the literature, their implementation has been prohibitive in general, except for linear models. We develop regularity and existence results on optimal policies. We then obtain rigorous approximation and near optimality results for jointly optimal coding and control policies. To this end, we first develop existence, regularity, and structural properties on optimal policies, followed by rigorous approximations and reinforcement learning results. Notably, we establish near optimality of finite model approximations obtained via predictor quantization as well as sliding finite window approximations, and their reinforcement learning convergence to near optimality. A detailed comparison of the approximation schemes and their reinforcement learning performance is presented.
I Introduction
Networked control systems involve stochastic control systems where a communication channel is present between different stations, such as sensors, actuators and controllers. In this context there exists a comprehensive literature on stabilization and optimization of such systems under various information constraints, see e.g. [1, 2, 3, 4, 5] for extensive reviews.
Networked control theory requires interdisciplinary methods to arrive at optimality and structural results on optimal coders, decoders and controllers. While analytically and architecturally very useful, the structural results on optimal policies often lead to optimization problems in uncountable spaces and hence lead to computationally challenging problems. Furthermore, despite the presence of such structural results in the literature as we will review, learning theoretic and rigorously justified near optimality results have been very limited. The goal of this paper is to present rigorous reinforcement learning algorithms which are provably near optimal for the jointly optimal coding and control policies in a system which is connected to a controller over a finite rate noiseless channel.
I-A Problem Setup
We consider a networked control problem where a controlled Markov source observed over a noiseless communication channel is controlled using data obtained from this channel. The controlled Markov source is updated at each time step with the following dynamics. For $t\geq 0,$
$x_{t+1}=f(x_{t},u_{t},w_{t}),$ (1)
where $x_{t}$ takes values from a finite state space $\mathds{X}$ [^1] The case with continuous spaces will be discussed in Section VII-A, $u_{t}$ takes values from a finite action space $\mathds{U}$, and $w_{t}$ is some independent and identically distributed (i.i.d.) noise process. The initial state $x_{0}$ is a random variable with distribution $\pi_{0}$.
At each time stage $t$, $x_{t}$ is encoded causally over a noiseless channel. The controller receives information from the communication channel and chooses a control action $u_{t}$ which is then transmitted to the plant. The encoder and controller use coding and control policies, respectively, to compute their outputs. This system is depicted in Figure 1.
Figure 1: Control driven Markov process over Noiseless channel
A coding policy $\gamma^{e}=\{\gamma^{e}_{t},t\geq 0\}$ is a sequence of functions which generate quantization outputs, $q_{t}$, as a measurable function of the encoder’s information at time $t$:
$I_{t}^{e}=\{{x_{[0,t]},q_{[0,t-1]}}\}$ (2)
$\gamma^{e}_{t}:I_{t}^{e}\mapsto q_{t}\in\mathcal{M}$ (3)
where $\mathcal{M}$ is a finite quantization output alphabet, so that $q_{t}=\gamma^{e}_{t}(I_{t}^{e})$ for $t\in\mathbb{Z}_{+}$.
A control policy $\gamma^{c}=\{\gamma^{c}_{t},t\geq 0\}$ is a sequence of functions which generate control actions, $u_{t}$, as a measurable function of the information available at the controller at time $t$:
$I_{t}^{c}=\{{q_{[0,t]},u_{[0,t-1]}}\}$ (4)
$\gamma^{c}_{t}:I_{t}^{c}\mapsto u_{t}\in\mathds{U}$ (5)
where $\mathds{U}$ is the finite action space, so that $u_{t}=\gamma^{c}_{t}(I_{t}^{c})$ for all $t\in\mathbb{Z}_{+}$.
Definition I.1
With the coding and control policies as given above, we define the class of admissible joint coding-control policies as
$\Gamma_{A}\coloneqq\{\overline{\gamma}=\{(\gamma^{e}_{t},\gamma^{c}_{t}),t\geq 0\}\}.$ (6)
The networked control problem involves a joint optimization of coding and control policies. The optimization objective for a finite time horizon is defined as follows for some $N\geq 0$:
$J^{N}(\pi_{0})=\inf_{\overline{\gamma}\in\Gamma_{A}}J^{N}(\pi_{0},\overline{\gamma})\coloneqq\inf_{\overline{\gamma}\in\Gamma_{A}}E_{\pi_{0}}^{\overline{\gamma}}\left[\sum^{N-1}_{k=0}c(x_{k},u_{k})\right].$ (7)
Here we use $E^{\overline{\gamma}}_{\pi_{0}}$ to denote the expectation on $(x_{t},u_{t})_{t\geq 0}$ given initial distribution $\pi_{0}$ and joint coding-control policy $\overline{\gamma}$. Our focus will be on an infinite horizon discounted cost criterion to be presented further below (19).
I-B Literature review and contributions
Before proceeding further, we would like to acknowledge some further related work, primarily in the control-free domain. In the control-free setup, related papers on real-time coding include the following: [6] established that the optimal causal encoder minimizing the data rate subject to a distortion constraint for an i.i.d. sequence is memoryless. If the source is $k$th-order Markov, then the optimal causal fixed-rate coder minimizing any measurable distortion uses only the last $k$ source symbols, together with the current state at the receiver’s memory [7]. Reference [8] considered the optimal causal coding problem of finite-state Markov sources over noisy channels with feedback. [9], and [10] considered optimal causal coding of Markov sources over noisy channels without feedback. [11] considered the optimal causal coding over a noisy channel with noisy feedback. Reference [12] considered the causal coding of more general sources, stationary sources, under a high-rate assumption. An earlier reference on quantizer design is [13]. Relevant discussions on optimal quantization, randomized decisions, and optimal quantizer design can be found in [14] and [15]. [16] have studied a related problem of coding of a partially observed Markov source. [17] considered within a multi-terminal setup decentralized coding of correlated sources when the encoders observe conditionally independent messages given a finitely valued random variable, and obtained separation results for the optimal encoders.
Further related studies include sequential decentralized hypothesis testing problems [18] and multi-access communications with feedback [19]. A detailed review is available in [4, Chapter 15]. Finally, [20] and [21] studied similar approximation and reinforcement learning techniques to those studied in this paper, but in the control-free case; the controlled setup requires a different and a more general MDP formulation and accordingly the analysis is significantly different.
Quantizer design for the linear quadratic Gaussian setup has been studied by many [22, 13, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Information theoretic methods lead to further methods, though with operational restrictions compared with zero-delay schemes: [33, 26, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43].
Contributions.
  • While structural results on the optimal encoding and control have been obtained in the literature especially for finite horizon problems, their implementation has been prohibitive in general, except for linear models. In this paper, we develop several regularity and structural properties leading to optimality results for jointly optimal coding and control.
  • We establish regularity properites such as weak Feller continuity, and filter/predictor stability. These ensure that the dynamic programming recursions are well defined and an optimal solutions exist under both finite and infinite horizon discounted optimization criteria.
  • These regularity results lead to rigorous near optimal approximation results where approximations are obtained either with quantized approximations or are obtained by finite window control policies, leading to near optimal solutions under complementary conditions.
  • Accordingly, both existence and near optimality results are new contributions to the literature.
  • Finally, corresponding to both of the approximation methods, we obtain rigorous reinforcement learning results obtained via associated quantized Q-learning algorithms and its convergence to near optimality: these involve reinforcement learning where the variables used in the algorithm are either the quantization of the predictor variables or are sliding finite window of the most recent measurement and action variables. Both of the algorithms are shown to converge to near optimality under complementary conditions. However, we present the comparative benefits of either of the methods in Section V.
I-C Notation
[…]
II Structural Results and Equivalent MDP Formulation
II-A Optimal Structure of Encoding and Control Policies
Define the predictor and filter sequences, respectively as
$\pi_{t}(\cdot):=P^{\overline{\gamma}}(x_{t}=\cdot\,|q_{[0,t-1]})$ (8)
$\overline{\pi}_{t}(\cdot)=P^{\overline{\gamma}}(x_{t}=\cdot\,|q_{[0,t]}).$ (9)
A celebrated structural result of Witsenhausen on the structure of optimal encoders [7] is extended in [31] and [4] to situations driven by control:
Theorem II.1
[7], [31], [4, Theorem 15.3.6]
For a system with dynamics (1), and optimization problem (7), any joint coding-control policy (with a given control policy) can be replaced, without loss in performance, by one which uses $x_{t}$ and $q_{[0,t-1]}$ (while not altering the control policy), at each $t\geq 0$. Equivalently,
$Q_{t}$ $=\gamma^{e}_{t}(q_{[0,t-1]}),\quad q_{t}=Q_{t}(x_{t}),$
where $Q_{t}:\mathbb{X}\to\mathcal{M}$ is a quantizer.
That is, the coding policy is a sequence of functions mapping the past encoder outputs to a finite space of all possible quantizers. The resulting quantizer, $Q_{t}$, then generates the quantization output $q_{t}\in\mathcal{M}$.
Walrand and Varaiya further refined Witsenhausen’s structural results in [8, 44], which include extensions to controlled setups (see also [31] and [4, Theorem 15.3.6]). The result shows that any control policy can, without loss in performance, be replaced by one that uses only $\pi_{t}$ and $q_{t}$ instead of all the information at the receiver. The version below follows from [4, Theorem 15.3.6] and its proof.
Theorem II.2
[44, Theorem 3.2][4, Theorem 15.3.6]
For a system with dynamics (1) and optimization problem (7), for any joint coding-control policy, the coding policy can be replaced, without loss in performance, by one which uses only $\pi_{t}$ and $x_{t}$, at each $t\geq 0$. Furthermore, the control policy can be replaced by one which uses only $\overline{\pi}_{t}$. That is, without any loss,
$Q_{t}$ $=\gamma^{e}_{t}(\pi_{t}),\quad q_{t}=Q_{t}(x_{t}),$
where as in the previous theorem $Q_{t}$ is a quantizer, and
$u_{t}=\gamma^{c}_{t}(\overline{\pi}_{t})\equiv\gamma^{c}_{t}(\pi_{t},Q_{t},q_{t}).$
In view of the above, we will define the following class of joint coder-control policies.
Definition II.1
A joint coder-controller policy $\overline{\gamma}=\{\gamma^{e},\gamma^{c}\}$, has a Controlled-Predictor-Structure if for each $t\geq 0$:
$\gamma^{e}_{t}:(\pi_{t},x_{t})\mapsto q_{t},\$ (10)
or equivalently,
$\gamma^{e}_{t}:\pi_{t}\mapsto{\cal Q}:=\{(Q_{t}:x_{t}\mapsto q_{t})\}.$ (11)
And, for each $t\geq 0$:
$\gamma^{c}_{t}:(\pi_{t},Q_{t},q_{t})\mapsto u_{t},$ (12)
or equivalently,
$\gamma^{c}_{t}:\pi_{t}\mapsto\mathcal{H}:=\Bigl{(}\eta_{t}:(Q_{t},q_{t})\mapsto u_{t}\Bigr{)}.$ (13)
where define :$\mathcal{H}:\{\mathcal{Q}\times\mathcal{M}\mapsto\mathcal{U}\}$ as the (finite) space of all $\eta_{t}$. We will denote this set of policies as $\Gamma_{C-P}$.
Accordingly, a policy in $\Gamma_{C-P}$ will map $\pi_{t}$ to $(Q_{t},\eta_{t})$ at time $t$. These are used to generate channel inputs and actions via $q_{t}=Q_{t}(x_{t})$ and $u_{t}=\eta_{t}(Q_{t},q_{t})$.
Thus, the effective state, as will be justified later in Theorem II.5, is $\pi_{t}$ and the action is $(Q_{t},\eta_{t})$.
Under a policy which satisfies the structure given in Theorem II.2, via a Bayesian update, we have the following predictor update equation:
$\pi_{t+1}(x_{t+1})$ (14)
$=\frac{\sum_{x_{t}}\sum_{u_{t}}P(x_{t+1}|x_{t},u_{t})P(q_{t}|\pi_{t},x_{t})P(u_{t}|q_{t},\pi_{t})\pi_{t}(x_{t})}{\sum_{x_{t+1}}\sum_{x_{t}}\sum_{u_{t}}P(x_{t+1}|x_{t},u_{t})P(q_{t}|\pi_{t},x_{t})P(u_{t}|q_{t},\pi_{t})\pi_{t}(x_{t})}$
Consider
$Q_{t}^{-1}(q_{t}):=\{x_{t}\in\mathbb{X};Q_{t}(x_{t})=q_{t}\}$
Then, we can write the update equation as
$\pi_{t+1}(x_{t+1})$ $=\frac{1}{\pi_{t}(Q_{t}^{-1}(q_{t}))}\sum_{Q_{t}^{-1}(q_{t})}P(x_{t+1}|x_{t},\eta_{t}(Q_{t},q_{t}))\pi_{t}(x_{t})$
$:=F(\pi_{t},Q_{t},\eta_{t},q_{t})$
II-B Predictor Structured Controlled Markov Problem
The structures for policies in $\Gamma_{C-P}$ motivate the following theorem, extended from [4, Theorem 15.4.1]. Throughout, we assume that we are using such a policy (which is without loss of optimality by Theorem II.2). Let $\mathcal{P}(\mathds{X})$ be the set of probability measures on $\mathds{X}$ endowed with weak convergence topology.
Theorem II.3
$(\pi_{t},(Q_{t},\eta_{t}))$ is a controlled Markov chain with $\pi_{t}$ (defined on $\mathcal{P}(\mathds{X})$) as the state and $(Q_{t},\eta_{t})$ (defined on $\mathcal{Q}\times\mathcal{H}$) as the control action.
Proof. Let $D\in\cal B({\cal P}(\mathbb{X}))$. Then, under a policy in $\Gamma_{C-P}$,
$P$ $(\pi_{t+1}\in D|\pi_{s},Q_{s},\eta_{s},s\leq t)$
$=\sum_{q_{t}\in\mathcal{M}}P(\pi_{t+1}\in D,q_{t}|\pi_{s},Q_{s},\eta_{s},s\leq t)$
$=\sum_{q_{t}\in\mathcal{M}}P(\pi_{t+1}\in D|q_{t},\pi_{s},Q_{s},\eta_{s},s\leq t)P(q_{t}|\pi_{s},Q_{s},\eta_{s},s\leq t)$
$=\sum_{q_{t}\in\mathcal{M}}P(F(\pi_{t},Q_{t},\eta_{t},q_{t})\in D|\pi_{t},Q_{t},\eta_{t},q_{t})P(q_{t}|\pi_{t},Q_{t},\eta_{t})$
$=P(F(\pi_{t},Q_{t},\eta_{t},q_{t})\in D|\pi_{t},Q_{t},\eta_{t})=P(\pi_{t+1}\in D|\pi_{t},Q_{t},\eta_{t})$
The third equality is due to the fact that:
$P($ $q_{t}|\pi_{s},Q_{s},\eta_{s},s\leq t)$
$=\sum_{\mathds{X}}P(q_{t},x_{t}|\pi_{s},Q_{s},\eta_{s},s\leq t)$
$=\sum_{\mathds{X}}P(q_{t}|x_{t},\pi_{s},Q_{s},\eta_{s},s\leq t)P(x_{t}|\pi_{s},Q_{s},\eta_{s},s\leq t)$
$=\sum_{\mathds{X}}P(q_{t}|x_{t},Q_{t})\tilde{P}(x_{t}|\pi_{t},Q_{t},\eta_{t})=P(q_{t}|\pi_{t},Q_{t},\eta_{t}),$
where we note that $q_{t}$ is determined by $x_{t}$ and $Q_{t}$ and since under $\overline{\gamma}\in\Gamma_{C-P}$, conditioning on $\pi_{t}$ implicitly conditions on $Q_{[0,t-1]}$ and $\eta_{[0,t-1]}$. $\diamond$
Thus we have an MDP with transition probability $P(d\pi_{t+1}|\pi_{t},Q_{t},\eta_{t})$. In order to apply approximation and learning algorithms to this redefined MDP, we build on [45, Chapter 4] and show that that the transition kernel is weakly continuous [46] (that is, the kernel is weak Feller).
Theorem II.4
The transition kernel $P(d\pi_{t+1}|\pi_{t},Q_{t},\eta_{t})$ has the weak Feller property, as defined in [47, C.3]. That is, it is weakly continuous in $\cal{P}(\mathds{X})\times\cal{Q}\times\cal{H}$ in the sense that for any $f\in C_{b}(\mathcal{P(\mathds{X})})$,
$\int_{\cal{P}(\mathds{X})}f(\pi_{t+1})P(d\pi_{t+1}|\pi_{t},Q_{t},\eta_{t})\in C_{b}(\mathcal{P(\mathds{X})}\times\mathcal{Q}\times\mathcal{H}).$
Proof. Fix some $t\geq 0$ and let $(Q^{n}_{t})_{n\geq 0}$ and $(\eta^{n}_{t})_{n\geq 0}$ be two sequences such that $Q^{n}_{t}\rightarrow Q_{t}$ and $\eta^{n}_{t}\rightarrow\eta_{t}$. Since $\cal{Q}$ and $\cal{H}$ are finite, this means that there exist some $N$ and $M$ such that for all $n\geq N$, $Q_{t}^{n}=Q_{t}$ and for all $n\geq M$, $\eta_{t}^{n}=\eta_{t}$. Thus, for a function $f$ on $C_{b}(\cal{P}(\mathds{X}))$ we have that for $n\geq\max(N,M)$,
$E[f(\pi_{t+1})|\pi_{t}^{n},Q^{n}_{t},\eta^{n}_{t}]=E[f(\pi_{t+1})|\pi_{t}^{n},Q_{t},\eta_{t}].$
With this, the proof then closely follows [46, Lemma 6 and Lemma 11]. Consider now:
$E[f(\pi_{t+1})|\pi_{t},Q_{t},\eta_{t}]$
$=\sum_{\cal{M}}f(\pi_{t+1})P(q_{t}|\pi_{t},Q_{t},\eta_{t})$
$=\sum_{\cal{M}}f(\pi_{t+1})\pi_{t}(Q_{t}^{-1}(q_{t}))$
$=\sum_{\cal{M}}f(F(\pi_{t},Q_{t},\eta_{t},q_{t}))\pi_{t}(Q_{t}^{-1}(q_{t})).$
Let $\pi_{t}^{n}\rightarrow\pi_{t}$ weakly. That is, for all continuous and bounded $g$, we have:
[…]
Note that since $\mathds{X}$ is finite, this is equivalent to $\pi^{n}_{t}\to\pi_{t}$ in total variation, i.e.,
[…]
where the supremum is over all measurable functions bounded by $1$. Now, we have that
$\quad\>\Bigl{|}\sum_{\cal{M}}f(\pi_{t+1})\pi_{t}(Q_{t}^{-1}(q_{t}))-\sum_{\cal{M}}f(\pi^{n}_{t+1})\pi^{n}_{t}(Q_{t}^{-1}(q_{t}))\Bigr{|}$
$\leq\Bigl{|}\sum_{\cal{M}}f(\pi_{t+1})\pi_{t}(Q_{t}^{-1}(q_{t}))-f(\pi_{t+1})\pi^{n}_{t}(Q_{t}^{-1}(q_{t}))\Bigr{|}$
$\qquad+\Bigl{|}\sum_{\cal{M}}f(\pi_{t+1})\pi^{n}_{t}(Q_{t}^{-1}(q_{t}))-f(\pi^{n}_{t+1})\pi^{n}_{t}(Q_{t}^{-1}(q_{t}))\Bigr{|},$ (15)
where $\pi^{n}_{t+1}=F(\pi^{n}_{t},Q_{t},\eta_{t},q_{t})$. By the fact that $\pi_{t}^{n}\rightarrow\pi_{t}$ in total variation, we have that $\pi^{n}_{t}(Q^{-1}_{t}(q_{t}))\to\pi_{t}(Q^{-1}_{t}(q_{t}))$, and thus the first term goes to 00.
For the second term, we have that
$\pi^{n}_{t+1}(\cdot)$ $=F(\pi_{t}^{n},Q_{t},\eta_{t},q_{t})$
$=\frac{1}{\pi^{n}_{t}(Q_{t}^{-1}(q_{t}))}\sum_{Q_{t}^{-1}(q_{t})}P(\cdot|x_{t},\eta_{t}(Q_{t},q_{t}))\pi^{n}_{t}(x_{t}),$
which, by the total variation convergence of $\pi^{n}_{t}$, converges to
$\frac{1}{\pi_{t}(Q_{t}^{-1}(q_{t}))}\sum_{Q_{t}^{-1}(q_{t})}P(\cdot|x_{t},\eta_{t}(Q_{t},q_{t}))\pi_{t}(x_{t})=\pi_{t+1}(\cdot)$
in total variation for every $q_{t}$ such that $\pi_{t}(Q_{t}^{-1}(q_{t}))>0$. Since $f$ is continuous, this implies that the second term in (15) converges to 0, and the result follows. $\diamond$
Note that a version of this theorem (in the non-controlled case) was proven in [46, Lemma 11], but for more general sources. Similar arguments could be applied here (see [46, Lemmas 3, 6, 11]) but since we are only considering the finite source/action setup here, the proof is more direct.
II-C Cost Equivalence and Optimality of Controlled-Predictor-Structured Policies
By Theorem II.2, for any $\overline{\gamma}\in\Gamma_{A}$, there exists some $\overline{\gamma}^{*}\in\Gamma_{C-P}$ such that
$E^{\overline{\gamma}^{*}}_{\pi_{0}}\left[\sum^{N-1}_{k=0}c(x_{k},u_{k})\right]\leq E^{\overline{\gamma}}_{\pi_{0}}\left[\sum^{N-1}_{k=0}c(x_{k},u_{k})\right].$
Now we have that
$E^{\overline{\gamma}^{*}}_{\pi_{0}}\left[\sum^{N-1}_{k=0}c(x_{k},u_{k})\right]=E^{\overline{\gamma}^{*}}_{\pi_{0}}\left[\sum^{N-1}_{k=0}E^{\overline{\gamma}^{*}}_{\pi_{0}}\left[c(x_{k},u_{k})\mid q_{[0,k-1]},Q_{k},\eta_{k}\right]\right]$
$=$ $E^{\overline{\gamma}^{*}}_{\pi_{0}}\left[\sum^{N-1}_{k=0}\sum_{\mathds{X}\times\mathds{U}\times\mathcal{M}}\eta_{k}(u_{k}|q_{k},Q_{k})Q_{k}(q_{k}|x_{k})\pi_{k}(x_{k})c(x_{k},u_{k})\right]$
$=$ $E^{\overline{\gamma}^{*}}_{\pi_{0}}\left[\sum^{N-1}_{k=0}\tilde{c}(\pi_{k},Q_{k},\eta_{k})\right],$
where
$\tilde{c}($ $\pi_{k},Q_{k},\eta_{k})$
$\coloneqq\sum_{\mathds{X}\times\mathds{U}\times\mathcal{M}}\eta_{k}(u_{k}|q_{k},Q_{k})Q_{k}(q_{k}|x_{k})\pi_{k}(x_{k})c(x_{k},u_{k})$ (16)
and the second equality follows from the fact that under $\overline{\gamma}^{*}\in\Gamma_{C-P}$, $Q_{k}$ and $\eta_{k}$ are deterministic functions of $\pi_{k}$ (and hence of $q_{[0,k-1]}$).
Using the Markov property of $(\pi_{k},Q_{k},\eta_{k})$ from Theorem II.3, we have thus a finite horizon cost criterion for a reformulated problem as
$J^{N}(\pi_{0},\overline{\gamma})=E^{\overline{\gamma}}_{\pi_{0}}\left[\sum^{N-1}_{k=0}\tilde{c}(\pi_{k},Q_{k},\eta_{k})\right]$ (17)
for some $\overline{\gamma}\in\Gamma_{C-P}$, where $\tilde{c}$ is the equivalent cost function at each effective state, $\pi_{t}$, and effective action, $(Q_{t},\eta_{t})$. Here we have written $\eta_{k}(u_{k}|q_{k},Q_{k})$ and $Q_{k}(q_{k}|x_{k})$ for simplicity of notation, noting that these deterministic functions can alternatively be viewed as conditional probabilities on their respective spaces. Further, let us define the minimum finite horizon cost criterion as
$J^{N}(\pi_{0}):=\inf_{\overline{\gamma}\in\Gamma_{C-P}}E^{\overline{\gamma}}_{\pi_{0}}\left[\sum^{N-1}_{k=0}\tilde{c}(\pi_{k},Q_{k},\eta_{k})\right].$ (18)
This is now a standard Markov Decision Problem which then admits an optimal policy (of Markov type) given the weak Feller condition (see Theorem II.4) under measurable selection conditions for weakly continuous kernels [47, Chapter 3].
Thus we have proven the following.
Theorem II.5
The minimum cost of the two problems (7) and (17) are equivalent. That is,
$\inf_{\overline{\gamma}\in\Gamma_{A}}E^{\overline{\gamma}}_{\pi_{0}}\left[\sum^{N-1}_{k=0}c(x_{k},u_{k})\right]=\inf_{\overline{\gamma}\in\Gamma_{C-P}}E^{\overline{\gamma}}_{\pi_{0}}\left[\sum^{N-1}_{k=0}\tilde{c}(\pi_{t},Q_{t},\eta_{t})\right]$
Accordingly, an optimal coding and control policy in $\Gamma_{C-P}$, when exists, is also optimal among all admissible coding and control policies.
II-D Infinite Horizon Discounted Cost Criterion
The infinite horizon discounted cost criteria for the reformulated problem is
$J_{\beta}(\pi_{0},\overline{\gamma})=\lim_{N\rightarrow\infty}J_{\beta}^{N}(\pi_{0},\overline{\gamma})$ (19)
where
$J_{\beta}^{N}(\pi_{0},\overline{\gamma})=E^{\overline{\gamma}}\left[\sum_{k=0}^{N-1}\beta^{k}\tilde{c}(\pi_{k},Q_{k},\eta_{k})\right],$
and $\beta\in(0,1)$ is the discount factor. The infinite horizon discounted cost problem is to find
$J_{\beta}(\pi_{0})=\inf_{\overline{\gamma}\in\Gamma_{A}}J_{\beta}(\pi_{0},\overline{\gamma})$ (20)
We have the following theorem, which is a standard result in the stochastic control literature under certain assumptions on an MDP (which we will show hold for our problem). Note that a stationary policy is one which does not depend on $t$ (i.e., $\gamma_{t}=\gamma$ for all $t\geq 0$).
Theorem II.6
For the system in Figure 1 and optimization objective (20), there exists an optimal policy in $\Gamma_{C-P}$ which is stationary.
Proof.
From Theorem II.5, we know that an optimal policy for (18) will be in $\Gamma_{C-P}$. Using the following inequality, we will now argue that a lower bound for (20) can be achieved by a minimizing stationary policy for (18).
$J_{\beta}(\pi_{0})=\inf_{\overline{\gamma}\in\Gamma_{A}}\lim_{N\rightarrow\infty}J_{\beta}^{N}(\pi_{0},\overline{\gamma})\geq\lim_{N\rightarrow\infty}\inf_{\overline{\gamma}\in\Gamma_{A}}J^{N}_{\beta}(\pi_{0},\overline{\gamma})$ (21)
First note that:
$J_{\beta}^{N}$ $(\pi_{0},\overline{\gamma})=$
$E_{\pi_{0}}^{\overline{\gamma}}[\tilde{c}(\pi_{0},Q_{0},\eta_{0})+\beta E[\sum_{k=1}^{N-1}\beta^{k-1}\tilde{c}(\pi_{k},Q_{k},\eta_{k})]|\pi_{0},Q_{0},\eta_{0}]$
Taking the infimum over $\overline{\gamma}\in\Gamma_{A}$ and noting that for finite horizons $\Gamma_{C-P}$ is an optimal class:
$J_{\beta}^{N}(\pi_{0})$ $=\inf_{\overline{\gamma}\in\Gamma_{A}}E_{\pi_{0}}^{\overline{\gamma}}[\tilde{c}(\pi_{0},Q_{0},\eta_{0})+\beta E[J_{\beta}^{N-1}(\pi_{1})|\pi_{0},Q_{0},\eta_{0}]$
$=\inf_{\overline{\gamma}\in\Gamma_{C-P}}E_{\pi_{0}}^{\overline{\gamma}}[\tilde{c}(\pi_{0},Q_{0},\eta_{0})+\beta E[J_{\beta}^{N-1}(\pi_{1})|\pi_{0},Q_{0},\eta_{0}]$
As $N$ increases, $J_{\beta}^{N}(\pi_{0})$ is also increasing monotonically and by a contraction argument [47] we have that $\lim_{N\rightarrow\infty}J_{\beta}^{N}(\pi_{0})=J^{\infty}_{\beta}(\pi_{0})$ exists,
$J^{\infty}_{\beta}(\pi_{0})\coloneq\lim_{N\rightarrow\infty}\min_{Q_{0},\eta_{0}}(\tilde{c}(\pi_{0},Q_{0},\eta_{0})+\beta E[J^{N-1}_{\beta}(\pi_{1})|\pi_{0},Q_{0},\eta_{0}])$
which thus serves as a lower bound to the optimal cost. The cost function $c(x,u)$ is bounded and since $\tilde{c}(\pi_{t},Q_{t},\eta_{t})$ is continuous in $\pi_{t}$, we have that $\tilde{c}(\pi_{t},Q_{t},\eta_{t})$ is continuous and bounded on $\mathcal{P}(\mathds{X})\times\mathcal{Q}\times\mathcal{H}$. Furthermore, $\cal{Q}\times\cal{H}$ is compact, and by Theorem II.4 we have that $P(d\pi_{t+1}|\pi_{t},Q_{t},\eta_{t})$ is weak Feller. Thus (see [48, Chapter 8.5]) there exists a stationary $\overline{\gamma}\in\Gamma_{C-P}$ which is optimal for (19) (that is, it satisfies (20)) (for a concise discussion, see [49, Lemma 5.5.4]). $\diamond$
II-E Interpretation of the Structural Results in Several Special Cases
The problem setup being considered is one which is dynamically controlled and in which the measurements are obtained via coding. This setup is a generalization of the following two setups:
  1. 1.
    Uncontrolled zero-delay coding setup. We will first consider the case where the system is not driven by control. Dynamics for this system are described by $x_{k+1}=f(x_{k},w_{k})$, for all $k\geq 0$. In this setup, the set of admissible encoder policies is
    $\gamma_{t}^{e}:I_{t}^{e}\to\cal{M},$
    where $I^{e}_{t}=(x_{[0,t]},q_{[0,t-1]})$. Instead of a controller we would have a decoder and a decoder policy, called $\gamma^{d}$. The decoder policy would map the channel output to a reconstruction of the source $\hat{x}_{t}$ defined on $\mathds{X}$. That is,
    $\gamma^{d}_{t}:{\cal{M}}^{t+1}\to\mathds{X}.$
    An optimization objective for this problem is to minimize the difference between the state and the reconstructed state. Let $\gamma^{e}=\{\gamma^{e}_{t}\}_{t\geq 0}$ and $\gamma^{d}=\{\gamma^{d}_{t}\}_{t\geq 0}$. The finite horizon cost is:
    $J^{N}(\pi_{0},\gamma^{e},\gamma^{d})=E_{\pi_{0}}^{\gamma^{e},\gamma^{d}}\Bigg{[}\sum^{N}_{t=0}d(x_{t},\hat{x}_{t})\Bigg{]}$
    where $d:\mathds{X}\times\hat{\mathds{X}}\mapsto\mathds{R}_{+}$ is a distance metric. The reconstructed state is not fed back into the plant and does not impact the next state value, $x_{t+1}$. We can view this problem as a special case of our framework by letting $u_{t}=\hat{x}_{t}$ and $x_{t+1}=f(x_{t},u_{t},w_{t})=f(x_{t},w_{t})$ for all $u_{t}$. Our “cost” would then become $c(x,u)=d(x,\hat{x})$.
    By Theorem II.2, we know that the optimal encoding-decoding policy is of the form $\pi_{t}\to(Q_{t},\eta_{t})$. In the non-controlled case, we can explicitly identify the optimal decoder for a given encoder by:
    $\gamma^{d}_{t}(q_{[0,t]})=\min_{\hat{x}}E_{\pi_{0}}^{\gamma^{e}}\left[d(x_{t},\hat{x})|q_{[0,t]}\right],$
    which indeed has the form $\gamma^{c}_{t}:\pi_{t}\mapsto\Bigl{(}\eta_{t}:(Q_{t},q_{t})\mapsto u_{t}\Bigr{)}$, as in (13). Similarly, it was previously shown in [8] that the optimal encoder has the form
    $\gamma^{e}_{t}:\pi_{t}\mapsto(Q_{t}:x_{t}\mapsto q_{t}),$
    as in (11).
    Since the decoder/controller can be explicitly identified in this case, and since the “control” doesn’t affect the source evolution, this can instead be studied as a Markov chain $(\pi_{t},Q_{t})$ to only search for the optimal encoder $\gamma^{e}$. This approach was used in [50, 51].
  2. 2.
    Partially Observed Markov Decision Problems (POMDPs) Instead of using an encoding policy, $\gamma^{e}_{t}$, to find a quantization of $x_{t}$, suppose that the quantization is fixed as $\tilde{Q}$. Such a fixed quantization induces a POMDP. The coding policy is a constant function:
    $\gamma^{e}_{t}:\pi_{t}\mapsto\tilde{Q}.$
    $\eta_{t}$ will now also be constant in $Q_{t}$. This is the same as a system with dynamics:
    $x_{t+1}$ $=f(x_{t},u_{t},w_{t})$
    $q_{t}$ $=\tilde{Q}(x_{t})$
    The controller will only have access to channel output $q_{t}$ making it analogous to an observation of $x_{t}$ at time t. When the control problem is structured in this form, the problem is reduced to that of a POMDP with observation $q_{t}$ of $x_{t}$. Structural results for POMDPs are well established [52, 53, 54] and it is well-known that an optimal control will be using the filter $\overline{\pi}_{t}$ as a sufficient statistic; here an alternative interpretation is provided which is consistent with the real-time encoding and control framework: the control policy will be structured as
    $\gamma^{c}_{t}:\pi_{t}\mapsto(\eta_{t}:q_{t}\mapsto u_{t})$
    or
    $\gamma^{c}_{t}:(\pi_{t},q_{t})\mapsto u_{t}.$
III Finite State Approximation via Predictor Quantization, and its Near Optimality and Reinforcement Learning
Due to the uncountable state space of probability measures on $\cal{P}(\mathds{X}$), computation of value functions is challenging. To address this problem, we will quantize this uncountable space to arrive at an approximate finite space model whose solution will be near-optimal for the original, building on the analysis introduced in [45, Chapter 4] for weakly continuous MDPs. Note that this quantization is different than the one that took place in Section I-A where a finite space $\mathds{X}$ was quantized into a smaller alphabet.
III-A Finite MDP Approximation and its Near Optimality
Quantization (or discretization) of an MDP on a possibly uncountable space $\mathds{Z}$ will be done as follows. We will define ${\mathds{Z}_{n}}=\{\hat{z}_{n,1},...,\hat{z}_{n,m_{n}}\}$ as a finite set that approximates $\mathds{Z}$ and a measurable mapping $\rho:\mathds{Z}\mapsto\mathds{Z}_{n}$. If we define a disjoint partition $\{B_{n,i}\}_{i=1}^{m_{n}}$ of $\mathds{Z}$, and pick a representative $\hat{z}_{n,i}$ in each $B_{n,i}$ such that
$\forall z\in B_{n,i},\rho(z)=\hat{z}_{n,i},$
then $\mathds{Z}_{n,i}$ can be considered a finite set of representative states for bounded sections of uncountable $\mathds{Z}$ [45]. More specifically, we will define $\rho(z)=\hat{z}$ as the nearest neighbour map:
$\hat{z}=\operatorname*{arg\,min}_{z^{\prime}\in\mathds{Z}_{n}}d_{z}(z,z^{\prime})$
where $d_{z}(\cdot,\cdot)$ is a metric on $\mathds{Z}$.
Building on [45, Chapter 4], we construct an approximate finite model and to establish near optimality of this approximate MDP we will follow [55, Section 2.3]. First, we will select a measure $\kappa\in\mathcal{P}(\mathds{Z})$, for which $\kappa(B_{n,i})\geq 0$ for all $B_{n,i}$. Using this measure, we can define a stage-wise cost and transition kernel. We will denote the cost and kernel from the original MDP as $c(z,u)$ and $P(dz^{\prime}|z,u)$, respectively. For $\hat{z}_{n,i},\hat{z}_{n,j}\in\mathds{Z}_{n}$ and $u\in\mathds{U}$ we define:
$c_{n}(\hat{z}_{n,i},u)\coloneqq\int_{B_{n,i}}\frac{\kappa(dz)}{\kappa(B_{n,i})}c(z,u)$ (22)
$P_{n}$ $(\hat{z}_{n,j}|\hat{z}_{n,i},u)\coloneqq\int_{z\in B_{n,i}}\int_{z^{\prime}\in B_{n,j}}P(dz^{\prime}\mid z,u)\frac{\kappa(dz)}{\kappa(B_{n,i})}$ (23)
We can now define our approximate MDP:
Definition III.1
For an MDP with state space $\mathds{Z}$, action space $\mathds{U}$, transition kernel $P(dz^{\prime}|z,u)$, and cost function $c(z,u)$, which we denote by $\text{MDP}=(\mathds{Z},\mathds{U},P(dz^{\prime}|z,u),c(z,u))$ we define the Finite State Approximate MDP as $\text{MDP}_{n}\coloneqq(\mathds{Z}_{n},\mathds{U},P_{n}(\hat{z}_{n,j}|\hat{z}_{n,i},u),c_{n}(\hat{z}_{n,i},u))$.
The optimal value function for $\text{MDP}_{n}$ is the solution
$\hat{J}_{\beta}:\mathds{Z}_{n}\mapsto\mathds{R}$
to the Discounted Cost Optimality Equation (DCOE):
$\hat{J}_{\beta}(\hat{z}_{n,i})=\min_{u\in\mathds{U}}(c_{n}(\hat{z}_{n,i},u)+\beta\sum_{\hat{z}_{n,j}\in\mathds{Z}_{n}}\hat{J}_{\beta}(\hat{z}_{n,j})P_{n}(\hat{z}_{n,j}\mid\hat{z}_{n,i},u)$
for $\hat{z}_{n,i}\in\mathds{Z}_{n}$. Note that we can also extend this function over $\mathds{Z}$ as follows: if $\hat{z}_{n,i}\in B_{n,i}$, then for all $z\in B_{n,i}$ we have $\hat{J}_{\beta}(z)\coloneqq\hat{J}_{\beta}(\hat{z}_{n,i})$ [45]. Equivalently, we can extend a policy $\hat{\gamma}_{n}$ for the $\text{MDP}_{n}$ to a policy $\tilde{\gamma}$ on the original MDP as follows.
$\tilde{\gamma}(z)=\hat{\gamma}_{n}(\hat{z})$ (24)
wherever $\hat{z}=\rho(z)$. We now state an assumption on the original MDP.
Assumption III.1
[55, Section 2.3][45, Chapter 4]
  1. (i)
    The cost function $c(z,u)$ is continuous and bounded.
  2. (ii)
    The transition kernel $P(dz^{\prime}|z,u)$ is weakly continuous.
  3. (iii)
    Spaces $\mathds{Z}$ and $\mathds{U}$ are compact.
Using the compactness of $\mathds{Z}$ from this assumption, we have that
$\lim_{n\rightarrow\infty}\max_{z\in\mathds{Z}}\min_{i=1,...,m_{z}}d_{z}(z,\hat{z}_{n,i})=0$
which motivates the following theorem.
Theorem III.1
[45, Theorem 4.3]
For all $z_{0}\in\mathds{Z}$ and $\tilde{\gamma}^{*}$, a policy extended over the MDP as in (24) from an optimal $\hat{\gamma}_{n}^{*}$ on $\text{MDP}_{n}$, we have
$\lim_{n\rightarrow\infty}|J_{\beta}(z_{0},\tilde{\gamma}^{*})-J_{\beta}(z_{0})|=0.$
That is, the optimal policy for $\text{MDP}_{n}$, when appropriately extended over $\mathds{Z}$, becomes near-optimal for the original MDP.
Consider now the Controlled-Predictor MDP introduced in Section II-B; that is, $\text{MDP}=(\mathcal{P}(\mathds{X}),\mathcal{Q}\times\mathcal{H},P(d\pi^{\prime}|\pi,Q,\eta),\tilde{c}(\pi,Q,\eta))$. We first define the quantization of the belief-space $\cal{P}(\mathds{X})$ to a finite $\mathcal{P}_{n}(\mathds{X})$, where $n$ represents the resolution of the quantization. Let $\mathcal{P}_{n}(\mathds{X})$ be the following finite set:
$\{\hat{\pi}\in\mathcal{P}(\mathds{X})\colon\hat{\pi}=\left[\frac{k_{1}}{n},...,\frac{k_{|\mathds{X}|}}{n}\right],k_{i}=0,...,n,i=1,...,|\mathds{X}|\}$
Using the nearest neighbour quantization specified above, we define $\rho$:
$\rho(\pi)\coloneq\operatorname*{arg\,min}_{\hat{\pi}\in\mathcal{P}_{n}(\mathds{X})}d(\pi,\hat{\pi})$
where $d(\cdot,\cdot)$ is a metric on $\mathcal{P}(\mathds{X})$ (say the $L^{1}$ distance, since $\mathds{X}$ is finite) [45]. The partition of $\mathcal{P}(\mathds{X})$ that is induced by $\rho$ is
$\{B_{n,i}\}=\{\pi\in\mathcal{P}(\mathds{X}):\rho(\pi)=\hat{\pi}_{n,i}\}.$
We define $c_{n}(\hat{\pi}_{n,i},Q,\eta)$ and $P_{n}(\hat{\pi}_{n,j}|\hat{\pi}_{n,i},Q,\eta)$ analogously to (22) and (23), and define $\text{MDP}_{n}\coloneqq(\mathcal{P}_{n}(\mathds{X}),\mathcal{Q}\times\mathcal{H},P_{n}(\hat{\pi}_{n,j}|\hat{\pi}_{n,i},Q,\eta),c_{n}(\hat{\pi}_{n,i},Q,\eta))$.
We know, from Theorem II.4, that the transition kernel $P(d\pi^{\prime}|\pi,Q,\eta)$ is weak Feller continuous. We have also discussed the boundedness and continuity of $\tilde{c}(\pi,Q,\eta)$ in the proof of Theorem II.6. Spaces $\cal{Q}$ and $\cal{H}$ are compact (finite). Since $\mathds{X}$ is finite, $\cal{P}(\mathds{X})$ is compact, so we have satisfied Assumption III.1. Therefore, from Theorem III.1, we have the following result concerning the optimality of the solution to $\text{MDP}_{n}$.
Theorem III.2
If the optimal policy $\hat{\gamma}^{*}_{n}$ for $\text{MDP}_{n}$ is extended to $\tilde{\gamma}_{n}^{*}$ on the original controlled-predictor MDP, we have for all $\pi_{0}\in\cal{P}(\mathds{X})$:
$\lim_{n\to\infty}\left|J_{\beta}(\pi_{0},\tilde{\gamma}_{n}^{*})-J_{\beta}(\pi_{0})\right|=0$ (25)
III-B Quantized Q-Learning
We will now consider a learning algorithm based on Watkins and Dayan’s Q-Learning algorithm [56] and its extension to Quantized Q-Learning [4]. Noting the above near optimality result, we will show that we can run the Q-learning algorithm on the approximate $\text{MDP}_{n}$ to achieve a near optimal result for the original. We will first comment on the convergence of Q-Learning for non-Markovian environments. The following is a result from [57]. Let us have the following sequences:
  1. 1.
    $\{y_{t}\}_{t\geq 0}$ defined on finite space $\mathds{Y}$.
  2. 2.
    $\{u_{t}\}_{t\geq 0}$ defined on finite space $\mathds{U}$.
  3. 3.
    $\{C_{t}\}_{t\geq 0}$ defined on $\mathds{R}$.
  4. 4.
    $\{\alpha_{t}\}_{t\geq 0}$, where $\alpha_{t}(y,u):\mathds{Y}\times\mathds{U}\rightarrow\mathds{R}$ is the learning rate at time $t$.
  5. 5.
    $\{\mathbf{Q}_{t}\}_{t\geq 0}$, where $\mathbf{Q}_{t}:\mathds{Y}\times\mathds{U}\rightarrow\mathds{R}$ and $\mathbf{Q}_{0}\equiv 0$, are the Q-factors at time $t$.
We define the following Q-learning iterations for every state-action pair $(y,u)\in\mathds{Y}\times\mathds{U}$:
$\mathbf{Q}_{t+1}(y,u)=$ $(1-\alpha_{t}(y,u))\mathbf{Q}_{t}(y,u)$
$+\alpha_{t}(y,u)(c_{t}+\beta\min_{u\in\mathds{U}}\mathbf{Q}(y_{t+1},u))$ (26)
Under the following assumption, we have a convergence theorem.
Assumption III.2
[57]
  1. (i)
    $\alpha_{t}(y,u)=\begin{cases}\frac{1}{1+\sum^{t}_{k=0}\mathds{1}_{\{y_{k}=y,u_{k}=u\}}}&\text{if }(y_{t},u_{t})=(y,u)\\ 0&\text{otherwise}\\ \end{cases}$
    and $\sum_{t=0}^{\infty}\alpha_{t}(y,u)=\infty$ almost surely.
  2. (ii)
    For some function $c^{*}:\mathds{Y}\times\mathds{U}\mapsto\mathds{R}$, we have that almost surely
    $\frac{\sum^{t}_{k=0}c_{k}\mathds{1}_{\{y_{k}=y,u_{k}=u\}}}{\sum^{t}_{k=0}\mathds{1}_{\{y_{k}=y,u_{k}=u\}}}\rightarrow c^{*}(y,u).$
  3. (iii)
    For any $f:\mathds{Y}\to\mathds{R}$, there is a measure $P^{*}$ such that almost surely
    $\frac{\sum^{t}_{k=0}f(y_{k+1})\mathds{1}_{\{y_{k}=y,u_{k}=u\}}}{\sum^{t}_{k=0}\mathds{1}_{\{y_{k}=y,u_{k}=u\}}}\rightarrow\int f(y^{\prime})P^{*}(dy^{\prime}|y,u).$
Theorem III.3
[57, Theorem 2.1].
  1. 1.
    $\mathbf{Q}_{t}(y,u)\rightarrow\mathbf{Q}^{*}(y,u)$ almost surely for each $(y,u)\in\mathds{Y}\times\mathds{U}$.
  2. 2.
    $\mathbf{Q}^{*}(y,u)$ is the solution to
    $\mathbf{Q}^{*}(y,u)=c^{*}(y,u)+\beta\sum_{y^{\prime}\in\mathds{Y}}\min_{u^{\prime}\in\mathds{U}}\mathbf{Q}^{*}(y^{\prime},u^{\prime})P^{*}(y^{\prime}|y,u)$
  3. 3.
    An optimal policy for $\text{MDP}_{n}=(\mathds{Y},\mathds{U},P^{*},c^{*})$ is given by: $\hat{\gamma}_{n}^{*}(y)\coloneq\operatorname*{arg\,min}_{u\in\mathds{U}}\mathbf{Q}^{*}(y,u)$.
In order to apply the above result to the MDP in the previous sections, we require the following assumption:
Assumption III.3
Let $(Q_{t},\eta_{t})$ be chosen uniformly from $\mathcal{Q}\times\mathcal{H}$ at each $t\geq 0$. Then the resulting Markov source process $\{x_{t}\}_{t\geq 0}$ is positive Harris recurrent (see e.g., [49, Definition 3.2.4]), and thus admits a unique invariant measure $\zeta$.
A sufficient condition for this assumption would be for the matrices $P(x^{\prime}|x,u)$ to be irreducible and aperiodic for each $u\in\mathds{U}$.
We apply the above algorithm to our (approximate) controlled-predictor MDP. Accordingly, let $y_{t}=\hat{\pi}_{t},u_{t}=(Q_{t},\eta_{t}),$ and $C_{t}=\tilde{c}(\pi_{t},Q_{t},\eta_{t})$. To show that this algorithm converges, we will need to satisfy Assumption III.2 for the sequence $(\hat{\pi}_{t},(Q_{t},\eta_{t}),\tilde{c}(\pi_{t},Q_{t},\eta_{t}),\alpha_{t})=(y_{t},u_{t},C_{t},\alpha_{t})$.
We will now present some results to validate these assumptions. We follow the strategy in [20, Section IV] to arrive at the unique ergodicity of the predictor process. Indeed, through a uniformly and independently chosen $Q_{t}$ and $\eta_{t}$, the system would be equivalent to a control-free hidden Markov model. Therefore by [20, Lemma 4] we have that the predictor process is stable in total variation and by [20, Theorem 4], we have unique ergodicity. Due to the specific nature of the problem, we also have the following direct argument for unique ergodicity.
First, we have that there exists some $Q\in\cal{Q}$, $x\in\mathds{X}$, $q\in\mathcal{M}$ such that $Q$ quantizes $x$ without any loss to $q$ (that is, $Q^{-1}(q)=\{x\}$). We also have, by the finiteness of $\cal{Q}$ and $\cal{H}$ and the positive Harris recurrence of $\{x_{t}\}_{t\geq 0}$ (under Assumption III.3), that for some $t\geq 0$ and all $\eta\in\cal{H}$, we have $(x_{t},Q_{t},\eta_{t})=(x,Q,\eta)$ almost surely [20]. This implies, through the update equation (14), that
$\pi_{t+1}(x^{\prime})$ $=\frac{1}{\pi_{t}(Q^{-1}(q))}\sum_{Q^{-1}(q)}P(x^{\prime}|x,\eta(Q,q))\pi_{t}(x)$
$=\frac{P(x^{\prime}|x,\eta(Q,q))\pi_{t}(x)}{\pi_{t}(x)}$
$=P(x^{\prime}|x,\eta(Q,q))$ (27)
Therefore, predictors of the form in (27) act as “atoms” with an almost surely finite return time, which implies that there is at most one invariant measure for $\{\pi_{t}\}_{t\geq 0}$ under the uniform selection of $(Q_{t},\eta_{t})$. Existence of an invariant measure is guaranteed by the compactness of $\mathcal{P}(\mathds{X})$ and the weak Feller property of $P(d\pi_{t+1}|\pi_{t},Q_{t},\eta_{t})$ (see [58, Theorem 7.2.3]). Thus, our process $\{\pi_{t}\}_{t\geq 0}$ admits a unique invariant measure, which we denote by $\lambda$. This leads to the following:
Lemma III.1
Assume that at each $k\geq 0$, $(Q_{k},\eta_{k})$ is chosen uniformly from $\mathcal{Q}\times\mathcal{H}$, and let Assumption III.3 hold. Then for any measurable and bounded function $f:\mathcal{P}(\mathds{X})\mapsto\mathds{R}$ and any $\pi_{0}\in P(\mathds{X})$,
[…]
Proof. This follows by the pathwise ergodic theorem [58, Corollary 2.5.2], and the fact that for any $\pi_{0}$, the hitting time to $\pi_{t}$ of the form in (27) is almost surely finite. $\diamond$
We will now define:
$B_{n}^{\lambda}\coloneqq\{B\in B_{n}:\lambda(B)>0\}$
and
$\mathcal{P}_{n}^{\lambda}(\mathds{X})\coloneqq\{\hat{\pi}\in\mathcal{P}_{n}(\mathds{X}):\rho^{-1}(\hat{\pi})\in B_{n}^{\lambda}).$
We then have the following:
Lemma III.2
[…]
  1. (i)
    $(\hat{\pi}_{t},Q_{t},\eta_{t})=(\hat{\pi},Q,\eta)$ infinitely often and $\sum_{t\geq 0}\alpha_{t}(\hat{\pi},Q,\eta)=\infty$.
  2. (ii)
    $\frac{\sum^{t}_{k=0}\tilde{c}(\pi_{k},Q_{k},\eta_{k})\mathds{1}_{\{\hat{\pi}_{k}=\hat{\pi},Q_{k}=Q,\eta_{k}=\eta\}}}{\sum^{t}_{k=0}\mathds{1}_{\{\hat{\pi}_{k}=\hat{\pi},Q_{k}=Q,\eta_{k}=\eta\}}}\rightarrow c_{n}(\hat{\pi},Q,\eta),$
    where
    $c_{n}(\hat{\pi},Q,\eta)\coloneqq\int_{B}\frac{\lambda(d\pi)}{\lambda(B)}\tilde{c}(\pi,Q,\eta)$
    and $B$ is the bin of $\hat{\pi}$.
  3. (iii)
    If $P_{n}(d\hat{\pi}^{\prime}|\hat{\pi},Q,\eta)$ is defined as in (23) using $\lambda$ instead of measure $\kappa$, we have
    $\frac{\sum^{t}_{k=0}f(\hat{\pi}_{k+1})\mathds{1}_{\{\hat{\pi}_{k}=\hat{\pi},Q_{k}=Q,\eta_{k}=\eta\}}}{\sum^{t}_{k=0}\mathds{1}_{\{\hat{\pi}_{k}=\hat{\pi},Q_{k}=Q,\eta_{k}=\eta\}}}\\ \to\int f(\hat{\pi}^{\prime})P_{n}(d\hat{\pi}^{\prime}|\hat{\pi},Q,\eta).$
Proof. Using Lemma III.1 with $(\pi_{t},Q_{t},\eta_{t})$, and noting that $(Q_{k},\eta_{k})$ are chosen uniformly, we have
$\frac{1}{t}\sum^{t-1}_{k=0}f(\pi_{k},Q_{k},\eta_{k})\rightarrow\int f(\pi,Q,\eta)\frac{\lambda}{|\cal{Q}\times\cal{H}|}(d\pi)$
as $t\rightarrow\infty$, almost surely. Defining
$f_{1}(\pi_{k},Q_{k},\eta_{k})\coloneqq\tilde{c}(\pi_{k},Q_{k},\eta_{k})\mathds{1}_{\{\hat{\pi}_{k}=\hat{\pi},Q_{k}=Q,\eta_{k}=\eta\}}$
and
$f_{2}(\pi_{k},Q_{k},\eta_{k})\coloneqq\mathds{1}_{\{\hat{\pi}_{k}=\hat{\pi},Q_{k}=Q,\eta_{k}=\eta\}}$
we then have
$\frac{\sum_{k=0}^{t}f_{1}(\pi_{k},Q_{k},\eta_{k})}{\sum_{k=0}^{t}f_{2}(\pi_{k},Q_{k},\eta_{k})}\to c_{n}(\hat{\pi},Q,\eta),$
which satisfies part (ii) of Assumptions III.2. Similarly, by letting $f_{1},f_{2}$ be the relevant functions, we also obtain parts (i) and (iii) (see also [21, Theorem 5.3]). $\diamond$
We can now state the following as a result of Theorem III.3 and as an extension of [20, Theorem 1].
Theorem III.4
[…]
$\mathbf{Q}^{*}($ $\hat{\pi},Q,\eta)=c_{n}(\hat{\pi},Q,\eta)$
$+\beta\sum_{\hat{\pi}^{\prime}\in\mathcal{P}_{n}^{\lambda}(\mathds{X})}P_{n}(\hat{\pi}^{\prime}\mid\hat{\pi},Q,\eta)\min_{(Q^{\prime},\eta^{\prime})\in\cal{Q}\times\cal{H}}\mathbf{Q}^{*}(\hat{\pi}^{\prime},Q^{\prime},\eta^{\prime})$
This coincides with the DCOE for $\text{MDP}_{n}=(\mathcal{P}^{\lambda}_{n}(\mathds{X}),\mathcal{Q},P_{n},c_{n})$, and therefore the policy $\hat{\gamma}^{*}_{n}(\hat{\pi})=\operatorname*{arg\,min}_{Q,\eta}\mathbf{Q}^{*}(\hat{\pi},Q,\eta)$ is optimal for $\text{MDP}_{n}$. Thus if we define a policy $\tilde{\gamma}_{n}^{*}(\pi)\coloneq\hat{\gamma}^{*}_{n}(\hat{\pi})$ where $\hat{\pi}=\rho(\pi)$, we have by Theorem III.2 that:
$\lim_{n\to\infty}|J_{\beta}(\pi_{0},\tilde{\gamma}_{n}^{*})-J_{\beta}(\pi_{0})|=0$
for any $\pi_{0}$ such that $\lambda(\pi_{0})>0$.
Remark: One such initialization where $\lambda(\pi_{0})>0$ would of course be those recurrent predictors identified in (27), or any $\pi$ which is reachable from one such $\pi_{0}$. Since $\pi_{t}$ has only finitely many possible values given $\pi_{0}$, each of these elements would also have positive measure under $\lambda$.
IV Sliding Finite Window Structured Policies, Their Near Optimality and Reinforcement Learning
Instead of the redefinition of the networked problem as its equivalent belief reduction, we will now consider another representation. We will use, as the state, a set of information which includes one previous belief term, and memory of a finite window of previous values of $Q_{t},q_{t},$ and $\eta_{t}$.
To introduce this structure, we will first define a window size $N$ and our new state sequence $\{\kappa_{t}\}_{t\geq 0}$, which is defined on $\mathds{W}=\mathcal{P}(\mathds{X})\times\mathcal{M}^{N}\times\cal{Q}^{N}\times\cal{H}^{N}$, by
$\kappa_{t}=\{\pi_{t-N},I^{N}_{t}\},$ (28)
where
$I_{t}^{N}=\{q_{[t-N,t-1]},Q_{[t-N,t-1]},\eta_{[t-N,t-1]}\}.$ (29)
$\pi_{t}$ can be found from $\kappa_{t}$ by starting from $\pi_{t-N}$ and using the update equation (14) $N$ times. This map will be denoted as $\psi:\mathds{W}\mapsto\mathcal{P}(\mathds{X})$, so that $\pi_{t}=\psi(\kappa_{t})$.
When at $t<N$, there will not be full window information. For simpler notation, at $t=0$, we will consider $\pi_{0}$ as found by updating $\pi_{-N}$ $N$ times.
Definition IV.1
We will say a joint coder controller policy has Controlled Finite Sliding Window Structure if, at time $t$, it uses only $\kappa_{t}$ to select $Q_{t}$ and $\eta_{t}$. We will denote this set of policies as $\Gamma_{C-FW}$.
As in section II-B, this motivates the definition of a new MDP.
Theorem IV.1
$(\kappa_{t},(Q_{t},\eta_{t}))$ is a controlled Markov chain on $\mathds{W}\times\cal{Q}\times\cal{H}$.
Proof.
$P(\kappa_{t+1}\in\cdot|\kappa_{s},Q_{s},\eta_{s},s\leq t)$
$=$ $P(\pi_{t-N+1},q_{[t-N+1,t]},Q_{[t-N+1,t]},\eta_{[t-N+1,t]}\in\cdot$
$\hskip 85.35826pt|\pi_{[0,t-N]},q_{[0,t-1]},Q_{[0,t]},\eta_{[0,t]})$
$=$ $P(\pi_{t-N+1},q_{[t-N+1,t]},Q_{[t-N+1,t]},\eta_{[t-N+1,t]}\in\cdot$
$\hskip 85.35826pt|\pi_{t-N},q_{[t-N,t-1]},Q_{[t-N,t]},\eta_{[t-N,t]})$
$=$ $P(\pi_{t-N+1},I_{t+1}\in\cdot|\pi_{t-N},I_{t},Q_{t},\eta_{t})$
$=$ $P(\kappa_{t+1}\in\cdot|\kappa_{t},Q_{t},\eta_{t}),$
where the second equality follows from the fact that $\pi_{t-N+1},q_{[t-N+1,t-1]},Q_{[t-N+1,t]}$ and $\eta_{[t-N+1,t]}$ are deterministic given $\pi_{t-N},q_{[t-N,t-1]},Q_{[t-N,t]},$ and $\eta_{[t-N,t]}$, and also that $Q_{t}$, $q_{t}$ depends on the past terms only through $\pi_{t}$ (which is a function of $\pi_{t-N},I_{t}^{N}$). $\diamond$
With an abuse of notation, define the following cost function:
$c(\kappa$ $,Q,\eta)=\sum_{(x,u,q)\in\mathds{X}\times\mathds{U}\times\mathcal{M}}\eta_{t}(u|Q,q)Q(q|x)\psi(\kappa)(x)c(x,u).$
Thus we obtain $\text{MDP}=(\mathds{W},\mathcal{Q}\times\mathcal{H},P(d\kappa^{\prime}|\kappa,Q,\eta),c(\kappa,Q,\eta))$. The objective is minimizing the following over all $\overline{\gamma}\in\Gamma_{A}$:
$J_{\beta}(\kappa_{0},\overline{\gamma})=E^{\overline{\gamma}}_{\kappa_{0}}\Big{[}\sum^{\infty}_{k=0}\beta^{k}c(\kappa_{k},Q_{k},\eta_{k})\Big{]}.$ (30)
We can extend the results of Theorem II.5 and Section II-D to argue that
$\inf_{\overline{\gamma}\in\Gamma_{A}}J_{\beta}(x_{0},\overline{\gamma})=\inf_{\overline{\gamma}^{\prime}\in\Gamma_{A}}J_{\beta}(\kappa_{0},\overline{\gamma}^{\prime})$
and also that an optimal policy for $J_{\beta}(\kappa_{0},\overline{\gamma})$ can be found in $\Gamma_{C-FW}$. That is,
$J_{\beta}(\kappa_{0})=\inf_{\overline{\gamma}\in\Gamma_{C-FW}}J_{\beta}(\kappa_{0},\overline{\gamma}).$
IV-A Near Optimality of Finite Sliding Window Approximation
We will fix the prior $\pi_{t-N}$ to be some (fixed) $\mu$. This approach follows that of [21, Sections 3.5] (see also [59]). Our state will now be $\hat{\kappa}_{t}=(\mu,I^{N}_{t})$ defined on $\hat{\mathds{W}}=\{\mu\}\times\cal{M}^{N}\times\cal{Q}^{N}\times\cal{H}^{N}$. In this section, we will consider the performance impact of using this approximation. Recalling the definition of $\pi_{t}$ in (8), we can define $\pi_{t}$ starting from the correct prior $\pi_{t-N}$ as:
$\pi_{t}(\cdot)=P^{\overline{\gamma}}_{\pi_{t-N}}(x_{t}=\cdot|q_{[t-N,t-1]},Q_{[t-N,t-1]},\eta_{[t-N,t-1]})$ (31)
and our approximate predictor that uses $\mu$ as the incorrect prior as:
$\hat{\pi}_{t}(\cdot)=P^{\overline{\gamma}}_{\mu}(x_{t}=\cdot|q_{[t-N,t-1]},Q_{[t-N,t-1]},\eta_{[t-N,t-1]}).$ (32)
We will define the following transition kernel for the finite sliding window approximation. Note that we now make the dependence on $N$ explicit as it affects the quality of the approximation (as will be shown shortly).
$P_{N}(\hat{\kappa}_{t+1}|\hat{\kappa}_{t},Q_{t},\eta_{t})$ $=P_{N}(\mu,I_{t+1}^{N}|\mu,I_{t}^{N},Q_{t},\eta_{t})$ (33)
$=P(\mathcal{P}(\mathds{X}),I^{N}_{t+1}|\mu,I^{N}_{t},Q_{t},\eta_{t}),$ (34)
where in the last line we are taking the marginal of the true transition kernel $P(d\kappa^{\prime}|\kappa,Q,\eta)$ over its first coordinate (the $\pi_{t-N}$ coordinate). We define the cost similarly:
$c_{N}(\hat{\kappa},Q,\eta)=\sum_{(x,u,q)\in\mathds{X}\times\mathds{U}\times\mathcal{M}}\eta(q,Q)Q(q|x)\psi(\hat{\kappa})(x)c(x,u).$ (35)
Then $\text{MDP}_{N}\coloneqq(\hat{\mathds{W}},\mathcal{Q}\times\mathcal{H},P_{N}(\hat{\kappa}^{\prime}|\kappa,Q,\eta),c_{N}(\hat{\kappa},Q,\eta))$. The discounted cost for $\text{MDP}_{N}$ will be $\hat{J}_{\beta}(\hat{\kappa}_{0},\hat{\gamma})$ and the optimal discounted cost is $\hat{J}_{\beta}(\hat{\kappa}_{0})$. By making $P_{N}$ and $c_{N}$ constant over all of $\cal{P}(\mathds{X})$ we will extend the optimal cost to $\tilde{J}_{\beta}(\kappa_{0})$. Let us say that $\hat{\gamma}^{*}_{N}$ achieves this optimal cost for the $\text{MDP}_{N}$ (which certainly exists here the space is finite [49]).
To study the performance of the approximation, we must consider the conditions that $\hat{\pi}_{t}$ can be used as an appropriate replacement for $\pi_{t}$. For a large $N$, this relies on filter stability which measures how quickly a process can recover from starting from the incorrect prior [21]. Let us define:
$L_{t}^{N}=\sup_{\gamma\in\Gamma_{C-FW}}E^{\gamma}_{\mu}[\|\pi_{t}-\hat{\pi}_{t}\|_{TV}]$ (36)
which measures the maximum total variation distance between the predictor with the correct and incorrect prior, at time $t$. From [60], we define the Dobrushin coefficient for $P\coloneqq P(x^{\prime}|x,u)$:
$\delta(P,u)=\min_{i,k\in\mathds{X}}\sum_{j\in\mathds{X}}\min(P(j|i,u),P(j|k,u)).$ (37)
We can now state the following, extending [61, Theorem 3.6], and recalling that $\overline{\pi}_{t}$ is the filter obtained by further conditioning $\pi_{t}$ on $q_{t}$. Also, to make the dependence on the initialization of $\pi_{0}$ clear, we write $\pi^{\mu}_{t}$ to denote the predictor process when $\pi_{0}=\mu$ (and similarly for the filter process).
Lemma IV.1
$\int_{\mathds{X}}$ $\int_{\mathcal{M}}\|\overline{\pi}_{t}^{\mu}-\overline{\pi}_{t}^{\nu}\|_{TV}P_{\mu}^{\gamma}(dq_{t}|x_{t},q_{[0,t-1]})\pi^{\mu}_{t}\leq 2\|\pi_{t}^{\mu}-\pi_{t}^{\nu}\|_{TV}.$
The proof of this follows from an identical process to the proofs of [21, Lemma 3]. This proof involves taking the Dobrushin coefficient of the conditional probability of the channel output given the state $x$ and the quantizer $Q$. We note that, since the channel is noiseless, the channel kernel $O(q^{\prime}|q)$, would be an identity map and $\delta(I)=0$. If the channel were noisy, the right side of this inequality would include a multiplier of $(2-\delta(O)$.
Theorem IV.2
For any $\mu\ll\nu$ and $\gamma\in\Gamma_{C-P}$, and if $\min_{u\in\mathds{U}}\delta(P,u)\geq 1/2,$
$E^{\gamma}_{\mu}[\|\pi_{t+1}-\hat{\pi}_{t+1}\|_{TV}]\leq 2(1-\min_{u\in\mathds{U}}\delta(P,u))E_{\mu}^{\gamma}[\|\pi_{t}-\hat{\pi}_{t}\|_{TV}]$ (38)
Proof. To prove this, we will show the statement for two instances of $\pi$ starting from distict priors $\mu$ and $\nu$, where $\mu\ll\nu$. Note that $\overline{\pi}_{t}^{\mu}$ can always be recursively computed when given $\mu$, $q_{[0,t]}$, under some policy $\gamma\in\Gamma_{C-P}$, so it is enough to only take the expectation with respect to $q_{[0,t]}$.
$E_{\mu}^{\gamma}[\|\overline{\pi}_{t}^{\mu}-\overline{\pi}_{t}^{\nu}\|_{TV}]$
$=\int_{\mathcal{M}^{t}}\int_{\mathcal{M}}\|\bar{\pi}_{t}^{\mu}-\bar{\pi}_{t}^{\mu}\|_{TV}P_{\mu}^{\gamma}(dq_{[0,t]})$
$=\int_{\mathcal{M}^{t}}\int_{\mathds{X}}\int_{\mathcal{M}}\|\bar{\pi}_{t}^{\mu}-\bar{\pi}_{t}^{\mu}\|_{TV}P_{\mu}^{\gamma}(dq_{t}|x_{t},q_{[0,t-1]})$
$\hskip 113.81102ptP_{\mu}^{\gamma}(dx_{t}|q_{[0,t-1]})P_{\mu}^{\gamma}(dq_{[0,t-1]})$
$\leq 2\int_{\mathcal{M}^{t}}\|\pi_{t}^{\mu}-\pi_{t}^{\nu}\|P_{\mu}^{\gamma}(dq_{[0,t-1]})$
$=2E^{\gamma}_{\mu}[\|\pi_{t}^{\mu}-\pi_{t}^{\nu}\|]$
The second last line is from Lemma IV.2. Since the Dobrushin coefficient is a contraction, we can state the following [60]:
$E_{\mu}^{\gamma}[\|\pi_{t+1}^{\mu}-\pi_{t+1}^{\nu}\|_{TV}]$
$\leq$ $(1-\min_{u\in\mathds{U}}\delta(P,u)E_{\mu}^{\gamma}[\|\overline{\pi}_{t}^{\mu}-\overline{\pi}_{t}^{\mu}\|_{TV}]$
$\leq$ $2(1-\min_{u\in\mathds{U}}\delta(P,u)E^{\gamma}_{\mu}[\|\pi^{\mu}_{t}-\pi_{t}^{\nu}\|_{TV}]$
$\diamond$
A bound for $L_{t}^{N}$ is now [21]:
$L_{t}^{N}$ $\leq\sup_{\gamma\in\Gamma_{C-P}}(2(1-\min_{u\in\mathds{U}}\delta(P,u)))^{N}E_{\mu}^{\gamma}[\|\pi_{t}-\hat{\pi}_{t}\|_{TV}]$
$\leq(2(1-\min_{u\in\mathds{U}}\delta(P,u)))^{N}\|\pi_{t}-\hat{\pi}_{t}\|_{TV}$
$\leq 4(1-\min_{u\in\mathds{U}}\delta(P,u)))^{N}$
We will use this loss bound to give a performance bound that shows the impact of using the finite window approximation. The following are extensions of [21, Theorems 1 and 2]. We will define $\|c\|_{\infty}\coloneq\max_{(x,u)\in\mathds{X}\times\mathds{U}}c(x,u)$.
Theorem IV.3
For any $\gamma\in\Gamma_{C-FW}$ which acts on the first $N$ time steps to generate $\kappa_{0}$ and any $\pi_{-N}\in\cal{P}(\mathds{X})$,
$E_{\pi_{-N}}^{\gamma}\big{[}|\tilde{J}_{\beta}(\kappa_{0})-J_{\beta}(\kappa_{0})]\leq\frac{2\|c\|_{\infty}}{(1-\beta)^{2}}(2(1-\min_{u\in\mathds{U}}\delta(P,u)))^{N}$
Proof.
We will show this for $N=1$. Note that $I_{t}^{1}=(q_{t},Q_{t},\eta_{t})$. By definition of $\tilde{J}_{\beta}$, we have $\tilde{J}_{\beta}(\kappa_{0})=\hat{J}_{\beta}(\hat{\kappa}_{0})$, and thus by the fixed-point equation for $\hat{J}_{\beta}$ (see e.g., [47, Chapter 4.2]), we have
$\sum_{q_{1}\in\mathcal{M}}$ $\tilde{J}_{\beta}(\pi_{0},(q_{0},Q_{0},\eta_{0}))P(q_{1}|\kappa_{0},Q_{0},\eta_{0})$
$=\sum_{q_{1}\in\mathcal{M}}\hat{J}_{\beta}(\mu,(q_{0},Q_{0},\eta_{0}))P(q_{1}|\kappa_{0},Q_{0},\eta_{0}).$
We add and subtract the above term, and use the fixed-point equations, to arrive at:
$E_{\pi_{-N}}^{\gamma}$ $\left[\left|J_{\beta}(\kappa_{0})-\tilde{J}_{\beta}(\kappa_{0})\right|\right]$
$\leq(\|c\|_{\infty}+\beta\|\hat{J}_{\beta}\|_{\infty})L_{0}^{1}+\sup_{\gamma^{\prime}\in\Gamma_{C-FW}}\beta E_{\pi_{-N}}^{\gamma^{\prime}}\left[\left|J_{\beta}(\kappa_{1})-\tilde{J}_{\beta}(\kappa_{1})\right|\right].$
where $L_{0}^{1}$ is defined as in (36). We apply the same process on the supremum term and recursively arrive at
$E_{\pi_{-N}}^{\gamma}\left[\left|J_{\beta}(\kappa_{0})-\tilde{J}_{\beta}(\kappa_{0})\right|\right]\leq\frac{\|c\|_{\infty}}{1-\beta}\sum^{\infty}_{k=0}\beta^{k}L_{k}^{1}$
where we used the fact that $\|J_{\beta}\|_{\infty}\leq\frac{\|c\|_{\infty}}{1-\beta}$. Using the results of Theorem IV.2 and the geometric series, we have the desired result:
$E_{\pi_{0}}^{\gamma}\left[\left|J_{\beta}(\kappa_{0})-\tilde{J}_{\beta}(\kappa_{0})\right|\right]$
$\leq\frac{\|c\|_{\infty}}{1-\beta}\sum^{\infty}_{k=0}\beta^{k}L^{1}_{k}$
$\leq\frac{\|c\|_{\infty}}{(1-\beta)^{2}}4(1-\min_{u\in\mathds{U}}\delta(P,u)))^{N}$
$\diamond$
Theorem IV.4
Let $\tilde{\gamma}_{N}^{*}$ be extended from $\hat{\gamma}^{*}_{N}$ over $\cal{P}(\mathds{X})$, where $\hat{\gamma}^{*}_{N}$ is optimal for $MDP_{N}$. Then for any $\gamma\in\Gamma_{C-P}$ which is applied $N$ times starting from $\pi_{-N}$ to generate $\kappa_{0}$, we have
$E_{\pi_{t-N}}^{\gamma}\big{[}J_{\beta}(\kappa_{0},\tilde{\gamma}_{N}^{*})-J_{\beta}(\kappa_{0})]\leq\frac{4\|c\|_{\infty}}{(1-\beta)^{2}}(2(1-\min_{u\in\mathds{U}}\delta(P,u)))^{N}$
Proof.
First we will state (again for $N=1$):
$E_{\pi_{t-N}}^{\gamma}\big{[}J_{\beta}(\kappa_{0},\tilde{\gamma}_{N}^{*})-J_{\beta}(\kappa_{0})]$
$\leq E_{\pi_{t-N}}^{\gamma}\big{[}J_{\beta}(\kappa_{0},\tilde{\gamma}_{N}^{*})-\tilde{J}_{\beta}(\kappa_{0})]$
$\hskip 113.81102pt+E_{\pi_{t-N}}^{\gamma}\big{[}\tilde{J}_{\beta}(\kappa_{0})-J_{\beta}(\kappa_{0})]$
Using a process similar to the one used to prove Theorem IV.3, we can get
$E_{\pi_{-N}}^{\gamma}\big{[}|J_{\beta}(\kappa_{0},\tilde{\gamma}^{*}_{N})-\tilde{J}_{\beta}(\kappa_{0})]\leq\frac{2\|c\|_{\infty}}{(1-\beta)^{2}}(2(1-\min_{u\in\mathds{U}}\delta(P,u)))^{N}$
and thus,
$E_{\pi_{t-N}}^{\gamma}\big{[}J_{\beta}(\kappa_{0},\tilde{\gamma}_{N}^{*})-J_{\beta}(\kappa_{0})]\leq\frac{4\|c\|_{\infty}}{(1-\beta)^{2}}(2(1-\min_{u\in\mathds{U}}\delta(P,u)))$
$\diamond$
IV-B Sliding Finite Window Q-Learning
Let us now consider the convergence of Q-learning to a solution for the finite sliding window $\text{MDP}_{N}$. Consider the following sequences:
  1. 1.
    $\{\hat{\kappa}_{t}\}_{t\geq 0}$ as defined above.
  2. 2.
    $\{Q_{t}\}_{t\geq 0}$ is chosen uniformly from $\cal{Q}$ at each time t.
  3. 3.
    $\{\eta_{t}\}_{t\geq 0}$ is chosen uniformly from $\cal{H}$ at each time t.
  4. 4.
    $\{C_{t}\}_{t\geq 0}$ where $C_{t}=c_{N}(\hat{\kappa}_{t},Q_{t},\eta_{t})$
  5. 5.
    $\{\alpha_{t}\}_{t\geq 0}$, where $\alpha_{t}:\hat{\mathds{W}}\times\cal{Q}\times\cal{H}\mapsto\mathds{R}_{+}$ is the learning rate at time $t$. Specifically, we define
    $\alpha_{t}(\hat{\kappa},Q,\eta)=\frac{1}{1+\sum^{t}_{k=0}\mathbf{1}_{\{(\hat{\kappa}_{k},Q_{k},\eta_{k})=(\hat{\kappa},Q,\eta)\}}}$
    if $(\hat{\kappa}_{t},Q_{t},\eta_{t})=(\hat{\kappa},Q,\eta)$, and 00 otherwise.
  6. 6.
    $\{\mathbf{Q}_{t}\}_{t\geq 0}$, where $\mathbf{Q}_{t}:\hat{\mathds{W}}\times\cal{Q}\times\cal{H}\mapsto\mathds{R}_{+}$, $\mathbf{Q}_{0}\equiv 0$, and $\mathbf{Q}_{t}$ is updated as:
    $\mathbf{Q}_{t+1}(\hat{\kappa},Q,\eta)=(1-\alpha_{t}(\hat{\kappa},Q,\eta))\mathbf{Q}_{t}(\hat{\kappa},Q,\eta)$
    $\hskip 28.45274pt+\alpha_{t}(\hat{\kappa},Q,\eta)[C_{t}+\beta\min_{(Q,\eta)\in\cal{Q}\times\cal{H}}\mathbf{Q_{t}}(\hat{\kappa}_{t+1},Q,\eta)]$
We will argue that the sequences $(\hat{\kappa}_{t},Q_{t},\eta_{t},C_{t},\alpha_{t})$ satisfy Assumption III.2, and thus the Q-Learning iterations will converge to a meaningful limit. As in the previous section, we require Assumption III.3, which we recall is that $\{x_{t}\}_{t\geq 0}$ is positive Harris recurrent (and thus has unique invariant measure $\zeta$) under the uniform choice of $(Q_{t},\eta_{t})$.
[…]
Lemma IV.2
[…]
  1. (i)
    $(\hat{\kappa}_{t},Q_{t},\eta_{t})=(\hat{\kappa},Q,\eta)$ infinitely often and $\sum_{t\geq 0}\alpha_{t}(\hat{\kappa},Q,\eta)=\infty$.
  2. (ii)
    $\frac{\sum^{t}_{k=0}C_{k}\mathds{1}_{\{\hat{\kappa}_{k}=\hat{\kappa},Q_{k}=Q,\eta_{k}=\eta\}}}{\sum^{t}_{k=0}\mathds{1}_{\{\hat{\kappa}_{k}=\hat{\kappa},Q_{k}=Q,\eta_{k}=\eta\}}}\rightarrow c_{N}(\hat{\kappa},Q,\eta)$
  3. (iii)
    For any $f$,
    $\frac{\sum^{t}_{k=0}f(\hat{\kappa}_{k+1})\mathds{1}_{\{\hat{\kappa}_{k}=\hat{\kappa},Q_{k}=Q,\eta_{k}=\eta\}}}{\sum^{t}_{k=0}\mathds{1}_{\{\hat{\kappa}_{k}=\hat{\kappa},Q_{k}=Q,\eta_{k}=\eta\}}}\\ \rightarrow\int_{\hat{\mathds{W}}}f(\hat{\kappa}^{\prime})P_{N}(d\hat{\kappa}^{\prime}|\hat{\kappa},Q,\eta)$
Proof. Under Assumption III.3, we have the following as $t\rightarrow\infty$: $\|P(x_{t}\in\cdot)-\zeta\|_{TV}\rightarrow 0$ as $t\rightarrow\infty$. Since for almost every $\hat{\kappa}$ we have that $P_{\zeta}(\hat{\kappa})>0$, we will eventually have $P_{\pi_{t-N}}(\hat{\kappa})>0$ and thus almost every $(\hat{\kappa},Q,\eta)$ is hit infinitely often, satisfying (i). Part (ii) is immediate because $C_{t}=c_{N}(\hat{\kappa},Q,\eta)$. The proof of part (iii) is as follows: because the marginals of $x_{t}$ converge to $\zeta$, under the random choice of $(Q_{t},\eta_{t})$, this causes the marginals on $I_{t}^{N}$ to also converge. Thus we have:
$\frac{\sum^{t}_{k=0}f(\hat{\kappa}_{k+1})\mathds{1}_{\{\hat{\kappa}_{k}=\hat{\kappa},Q_{k}=Q,\eta_{k}=\eta\}}}{\sum^{t}_{k=1}\mathds{1}_{\{\hat{\kappa}_{k}=\hat{\kappa},Q_{k}=Q,\eta_{k}=\eta\}}}$
$=\frac{\sum^{t}_{k=0}f(I_{k+1})\mathds{1}_{\{I_{k}=i,Q_{k}=Q,\eta_{k}=\eta\}}}{\sum^{t}_{k=1}\mathds{1}_{\{I_{k}=i,Q_{k}=Q,\eta_{k}=\eta\}}}$
$\rightarrow\int f(i^{\prime})P(\pi,i^{\prime}|\zeta,i,Q,\eta)$
$=\int_{\hat{\mathds{W}}}f(\hat{\kappa}^{\prime})P_{N}(d\hat{\kappa}^{\prime}|\hat{\kappa},Q,\eta)$
$\diamond$
We can now state the following as a result of Theorem III.3.
Theorem IV.5
If $\delta(P,u)>1/2$ for all $u\in\mathds{U}$, and under Assumption III.3, we have the following: for almost every $(\hat{\kappa},Q,\eta)\in\hat{\mathds{W}}\times\cal{Q}\times\cal{H}$, $\mathbf{Q}_{t}(\hat{\kappa},Q,\eta)$ will converge almost surely to a limit satisfying:
$\mathbf{Q}^{*}($ $\hat{\kappa},Q,\eta)=c_{N}(\hat{\kappa},Q,\eta)$
$+\beta\sum_{\hat{\kappa}^{\prime}\in\hat{\mathds{W}}}P_{N}(\hat{\kappa}^{\prime}\mid\hat{\kappa},Q,\eta)\min_{(Q^{\prime},\eta^{\prime})\in(\cal{Q}\times\cal{H})}\mathbf{Q}^{*}(\hat{\kappa^{\prime}},Q^{\prime},\eta^{\prime})$
This coincides with the DCOE for $\text{MDP}_{N}$, and therefore the policy $\hat{\gamma}^{*}_{N}(\hat{\kappa})=\operatorname*{arg\,min}_{Q,\eta\in{\cal{Q}\times\cal{H}}}\mathbf{Q}^{*}(\hat{\kappa},Q,\eta)$ is optimal for $\text{MDP}_{N}$. Thus if we define a policy $\tilde{\gamma}_{N}^{*}(\kappa)\coloneq\hat{\gamma}^{*}_{N}(\hat{\kappa})$, we have by Theorem IV.4 that for almost every $\kappa_{0}$,
$\lim_{N\to\infty}|J_{\beta}(\kappa_{0},\tilde{\gamma}_{N}^{*})-J_{\beta}(\kappa_{0})|=0.$
V Comparison of Finite Predictor State and Finite Window Approximations
We will now compare the two rigorously justified approximation methods presented: The finite sliding window method and quantized predictor approach.
  • ([i]) [Computational efficiency and availability of model information] The finite sliding window method is more computationally efficient because all possible values of $\hat{\kappa}_{t}$ and $\hat{\pi}_{t}=\psi(\hat{\kappa}_{t})$ can be computed offline before running the learning algorithm. However, the number of possible finite windows grows very fast in the relevant alphabet sizes. Because of this, each iteration requires less computation but, the number of Q-iterations before the algorithm converges to an optimal policy can be very large for window sizes larger than 2. Conversely, the computational complexity at each Q-iteration in the implementation of the quantized predictor approach is significantly higher. Learning requires frequent updates of the belief state (and requires a Bayesian update which necessitates access to the system model), and as the quantization becomes finer, the number of possible states grows exponentially (although not every state may be relevant, as we only need consider those with positive measure under the unique invariant measure, see Theorem III.4). Thus, even for resolutions of quantization up to 15, the number of visited states is small enough that the algorithm can converge relatively faster than the sliding window method (for window sizes greater than 2). This leads to a trade-off between approximation accuracy and computational costs for both these methods.
  • ([ii]) [Initialization] The sliding window method is insensitive to initialization. Provided the window length is sufficiently large, it converges to near-optimal policies regardless of the initial distribution, making it advantageous in real-world applications where the initial state might not be precisely known or controllable. The quantized predictor state space method is sensitive to initialization. The learned policies from this approach are near-optimal only when the initial state has positive measure under the unique invariant measure for $\pi_{t}$. This may be problematic in practical scenarios where the system may start from an arbitrary initial condition.
  • ([iii]) [On required filter stability] The finite sliding window method assumes a strong form of controlled filter stability. Only under appropriate conditions, such as those characterized by the Dobrushin coefficient, can this approach have near-optimal performance. The quantized state space method can work under weaker conditions of stability, and does not require an assumption on the Dobrushin coefficient.
  • ([iv]) Overall, both methods can achieve near-optimality under the right conditions, but they are more suitable in complementary settings: The sliding window approach is more computationally efficient and more robust to changes in initial conditions, making it a better choice for applications where faster operation and ease of implementation are critical. The quantized state space method, while computationally less efficient, offers flexibility and can be more accurate in scenarios where the source’s initial distribution is well-known and the computational resources to handle the Bayesian updates are available.
VI Simulations
In this section, we provide simulation studies where both of the approximation methods introduced in the paper are implemented in the following.
VI-A Simulation: Predictor Quantization Based Approach
We will now give an example of the control problem and simulate the performance of the algorithm using the finite state approximation. We will use a discount factor of $\beta=0.8$.
Let $\mathds{X}=\{1,2,3\}$, $\mathds{U}=\{1,2\}$, and =$\mathcal{M}=\{1,2\}$. The transition kernel $P(\cdot|x,u)$ is as follows:
$P(\cdot|x,1)=\frac{1}{10}\begin{pmatrix}4&0&6\\ 3&7&0\\ 2.5&2.5&5\\ \end{pmatrix}$
$P(\cdot|x,2)=\frac{1}{10}\begin{pmatrix}3.5&5&1.5\\ 5&5&0\\ 0&7.5&2.5\\ \end{pmatrix}$
The cost function is defined as $c(x,u)=C_{(xu)}$, that is the element at $[x,u]$ of the cost matrix:
$C=\begin{bmatrix}0&0\\ 0&1\\ 1&1\\ \end{bmatrix}$ (39)
By design, the kernel is aperiodic and irreducible. For a randomly and independently generated first action, $u_{0}$, the unique invariant distribution of $P(\cdot|x,u_{0})$ is computed and used as the initial distribution for learning the best policy. For estimating the cost under this policy, $\pi_{0}$ is a predictor of the form in (27). For quantization resolutions $n=\{1,...5\}$, the aforementioned Q-learning algorithm is applied and policies (after $\sim 10^{6}$ iterations or earlier convergence) are applied. We calculate the empirical expected discounted cost for the learned policy by running the simulation with a horizon $N=1000$ and averaging this cost over 1000 Monte-Carlo iterations. As the resolution of quantization increases, the cost decreases, as expected. The results are shown in Figure 2.
Figure 2: Discounted cost using a learned policy for $\text{MDP}_{n}$ for quantization resolutions $n=\{1,3,5,10,15\}$
VI-B Simulation: Sliding Window Approximation Method
To simulate the performance of the algorithm using the finite sliding window approximation, we will use the same $\beta$, $\mathds{X}$, $\mathds{U}$, $\mathcal{M}$, and $c(x,u)$, as in Section VI-A above. This example will use the following kernel.
$P(\cdot|x,1)=\frac{1}{10}\begin{pmatrix}4&0&6\\ 4.5&3&2.5\\ 2.5&2.5&5\\ \end{pmatrix}$
$P(\cdot|x,2)=\frac{1}{10}\begin{pmatrix}3.5&5&1.5\\ 5&5&0\\ 0.5&7.5&2\\ \end{pmatrix}$
This kernel is again irreducible and aperiodic but, has been designed to also satisfy the Dobrushin coefficient condition in Theorem IV.3. That is, we have that $\min_{u\in\mathds{U}}\delta(P,u)=0.55\geq 1/2$ which means the results of Theorems IV.2 and IV.3 are applicable.
The source begins from $x_{-N}=1$, and the first $N$ steps are obtained by using a uniform choice of $(Q_{t},\eta_{t})$. A learned policy for the $\text{MDP}_{N}$ for window lengths $N=\{1,2,3,4,5\}$ was found by training the Q-learning algorithm for $10^{5}$ iterations. Using this policy, the empirical expected cost was computed using $N=1000$ and 1000 Monte-Carlo iterations. As the window length increases, the discounted cost decreases consistently, as expected. The results are shown in Figure 3.
Figure 3: Discounted cost using a learned policy for $MDP_{N}$ for window lengths $N=\{1,2,3,4,5\}$.
VII Extensions
In this section, we comment on two generalizations of our work presented in the paper. While we do not study them in detail due to space constraints, we provide a detailed discussion on how the extensions can be made with little conceptual effort, though with some additional technical analysis as explicitly detailed in the following.
VII-A Case with real or Polish State $\mathbb{X}$ and Action $\mathbb{U}$ Spaces
Many engineering systems involve continuous vector spaces. While we leave a detailed study of this direction for future work, also in view of space constraints, in the following we highlight the program that the extension would entail.
The structural result we have developed applies also to the case with general, standard Borel, $\mathbb{X},\mathbb{U}$; as studied in [4, Chapter 15] building on [46].
Assumption VII.1
[46] Let a controlled system be expressed in the stochastic realization form
$x_{t+1}=f(x_{t},u_{t},w_{t}),\quad t=0,1,2,...,$ (40)
where $f:\mathbb{R}^{d}\times\mathbb{U}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a Borel measurable function and ${w_{t}}$ is an independent and identically distributed (i.i.d.) vector noise sequence which is independent of $x_{0}$. It is assumed that for each fixed $x,u\in\mathbb{R}^{d}\times\mathbb{U}$, the distribution of $f(x,u,w_{t})$ admits the (conditional) density function $\varphi(\cdot|x,u)$ (with respect to the $d$-dimensional Lebesgue measure) which is positive everywhere. Furthermore, $\varphi(\cdot|x,u)$ is bounded and Lipschitz uniformly in $x,u$.
Definition VII.1
Let $\mathcal{G}$ denote the set of all probability measures on $\mathbb{R}^{d}$ admitting densities that are bounded by a constant $C$ and Lipschitz with constant $C_{1}$.
Note that viewed as a class of densities, $\mathcal{G}$ is uniformly bounded. In [46, Lemma 3] it is shown that $\mathcal{G}$ is closed in $\mathcal{P}(\mathbb{R}^{d})$ under weak convergence.
As discussed in [46], a quantizer $Q$ with cells $\{B_{1},\ldots,B_{M}\}$ can be characterized as a stochastic kernel $Q$ from $\mathbb{R}^{n}$ to $\{1,\ldots,M\}$ defined by
$Q(i|x)=1_{\{x\in B_{i}\}},\quad i=1,\ldots,M.$
Let us endow the quantizers with the topology induced by a stochastic kernel interpretation under the Young topology [46]. If $P$ is a probability measure on $\mathbb{R}^{n}$ and $Q$ is a stochastic kernel from $\mathbb{R}^{n}$ to ${\cal M}$, then $PQ$ denotes the resulting joint probability measure on $\mathbb{R}^{n}\times{\cal M}$. That is, a quantizer sequence $Q_{n}$ converges to $Q$ weakly at $P$ ($Q_{n}\to Q$ weakly at $P$) if $PQ^{n}\to PQ$ weakly. Similarly, $Q_{n}$ converges to $Q$ in total variation at $P$ ($Q_{n}\to Q$ at $P$ in total variation at $P$) if $PQ^{n}\to PQ$ in total variation.
For compactness properties [62, 63] [15], we restrict the set of quantizers considered by only allowing quantizers having convex quantization bins (cells) $B_{i}$, $i=1,\ldots,M$.
Assumption VII.2
The quantizers have convex codecells with at most a given number of cells; that is the quantizers live in $\mathcal{Q}_{c}(M)$, the collection of $k$-cell quantizers with convex cells where $1\leq k\leq M$.
In this context, let $\Gamma_{C-P}$ denote the set of all predictor structured controlled Markov policies which in addition satisfy the condition that all quantizers $Q_{t}$, $t\geq 0$ have convex cells (i.e., $Q_{t}\in\mathcal{Q}_{c}(M)$ for all $t\geq 0$).
For this problem, one can then obtain the necessary weak Feller regularity on the resulting MDP kernel $P(d\pi_{t+1}|\pi_{t},Q_{t},\eta_{t})$, similarly to [46, Lemma 11]. Accordingly, an existence result, and therefore a counterpart to Theorem II.5, can be obtained.
To obtain finite approximations, using the weak Feller regularity, one can quantize the source space and obtain an approximate Markov model [64, Section 3], show that the approximate model is close to the original model under weak convergence, and therefore show that the solution of the approximate model is near-optimal for the original model [55].
Accordingly, the analysis in the current paper, by applying an optimal zero-delay coding and control for an approximate finite model, will be near-optimal for the original model under mild technical conditions. A detailed analysis of this practically significant setup is left for future work, with refinements on the technical conditions to be presented.
VII-B Case with Noisy Channels and Feedback
Consider the case in Figure 4 in which there is a noisy channel between the quantization output $q_{t}$ and the channel output $q^{\prime}_{t}$ where the channel is a Discrete Memoryless Channel (DMC), which satisfies the property that $P(q^{\prime}_{t}|q_{[0,t]})=P(q^{\prime}_{t}|q_{t})$ for all realizations and history.
Figure 4: Optimal coding with a single encoding terminal. We will study various scenarios, a general version being a controlled model over a noisy channel.
The controlled separation results of the type given in Theorem II.2 hold in the noisy channel case as well by [8] and [4, Theorem 15.3.8], provided that the encoder has access to the realizations $q^{\prime}_{t}$ in a causal fashion.
Accordingly, for such a setup, the structural, existence, and approximation results presented in this paper generalize with essentially identical arguments presented in the paper provided that there is feedback from the channel output to the encoder input, see the analysis [65, Section IV] and [21] for the control-free setup.
Accordingly, the counterparts of the results in Sections II-IV apply nearly identically. One technical detail (for the learning results in Section III) is that in the noisy channel case we do not necessarily have recurrent values of $\pi_{t}$ (of the form in (27)). Thus, additional filter stability conditions are needed to ensure the uniqueness of an invariant measure of $\pi_{t}$ for the exploration process (see [21] for details in the control-free case). Another minor difference is that the Dobrushin coefficient term in Theorem IV.3 (and the related supporting results) have to be modified to include the effect of the channel.
To summarize, we have the following remark.
Remark VII.1
We can state the following counterparts in the noisy channel setup:
  • (i)
    For the minimization of the infinite horizion discounted cost (19) in the noisy channel case, quantizing the resulting MDP leads to a near-optimal policy, i.e., a noisy channel analog of Theorem III.2 holds.
  • (ii)
    One can obtain an optimal policy for the quantized MDP in (i) through Q-learning, i.e., a noisy channel analog of Theorem III.4 holds.
  • (iii)
    For the minimization of the infinite horizion discounted cost (19) in the noisy channel case, using a finite memory of past channel outputs and controls leads to a near-optimal policy, i.e., a noisy channel analog of Theorem IV.3 holds.
  • (iv)
    One can obtain an optimal policy for the finite memory MDP in (iii) through Q-learning, i.e., a noisy channel analog of Theorem IV.5 holds.
VIII Conclusion
This work has addressed the problem of optimal control over a finite-rate noiseless communication channel, proving structural results that characterize optimal coding and control policies for a Markovian system. For the purpose of Q-learning, we introduced two methods to approximate this MDP: the finite sliding window and the quantized state space approaches. We demonstrated the effectiveness of Q-learning for both methods and noted the trade-offs between computational efficiency and initialization sensitivity. While the sliding window approach is more practical for real-time systems due to its lower complexity, the quantized state space method allows for finer approximations but at a higher computational cost. Future research could extend these methods to more complex scenarios, such as noisy channels, channels with feedback, and systems with infinite state and action spaces.